8. (15 points) A solenoid 10 cm in diameter and 630 cm long is made from copper wire of diameter 0.050 cm, with very thin insulation. The wire is wound onto a cardboard tube to make a solenoid. (a) (5 points) Determine the resistance in the wire. You may take the resistivity of copper to be 1.68 x 10-62 (b) (5 points) The number of turns in the wire is ten. Find the current necessary to produce a magnetic field of strength 1.5T (c) (5 points) How much power is needed to produce the magnetic field?
Answer #1

The equation for resistance is

R = rho L / A = (1.68 x10^(-8) (6.3) / pi (0.025x10^(-2))^2 = 0.54 ohms

b) Magnetic field is

B = mu_o N I / L

1.5 = ( 4 \pi x 10^(-7) ) (10) I / 6.3

I = 7.52 x10^5 A

c)

Power P = I^2 R

   = (7.52x10^5)^2 (0.54) = 3.05 x10^(11) A